If PQ-QP=27 and P+Q=11,then find the value of P and Q ?

Grade:10

1 Answers

Himani Belsare
6 years ago
If P = 7, Q= 4, then 7+4 = 11 (satisfies 2nd eq)
and
74 – 47 = 27 ( it satisfies 2nd eq as well)
 
Therfore the values of P & Q are 7 and 4 respectively..
 Cheers!!